A rectangular parking lot is 100 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 500 ft diagonally.
Answer by nerdybill(7384)

A rectangular parking lot is 100 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 500 ft diagonally.
.
Let x = width
x+100 = length
.
Applying Pythagorean theorem:
x^2 + (x+100)^2 = 500^2
x^2 + x^2 + 200x + 10000 = 250000
2x^2 + 200x - 240000 = 0
x^2 + 100x - 120000 = 0
.
Factoring we get:
(x-300)(x+400) = 0
x = {-400, 300}
Only the positive answer makes sense so
x = 300 feet (width)
.
x+100 = 400 feet (length)
.
Conclusion:
parking lot is 300 feet by 400 feet